# shapes, arrows.meta, positioning Chorus: Foundation Models for Unified Data Discovery and Exploration

### v1

@@ Abstract.
We explore the application of foundation models to data discovery and exploration tasks. Foundation models are large language models (llms) that show promising performance on a range of diverse tasks unrelated to their training. We show that these models are highly applicable to the data discovery and data exploration domain. When carefully used, they have superior capability on three representative tasks: table-class detection, column-type annotation and join-column prediction. On all three tasks, we show that a foundation-model-based approach outperforms the task-specific models and so the state of the art. Further, our approach often surpasses human-expert task performance. This suggests a future direction in which disparate data management tasks can be unified under foundation models.
@@ 1. Introduction
Despite their promise, serious risks have hampered the reception of foundation models. These include: spurious generation (including “hallucination”) (Holtzman et al., 2020), factual recall limitations (Mallen et al., 2022), bias (Gehman et al., 2020), dataset contamination (Dodge et al., 2021), logical shortcuts (Shi et al., 2023) and fallacies (Lin et al., 2022). Naïve deployment can lead to unanticipated problems: it has already led to legal action (Court, 2023) and recalls by major corporations (Heaven, 2022). These risks are now acknowledged by even the creators of these models (Bubeck et al., 2023; Ouyang et al., 2022; Touvron et al., 2023).
The goal of this paper is to demonstrate the utility of foundation models to the data discovery and exploration domain while mitigating the aforementioned risks. We select three representative tasks to show the promise of foundation models: 1⃝ table-class detection, 2⃝ column-type annotation and 3⃝ join-column prediction. An outline of our approach is shown in Figure 1. We call this approach chorus.
Preprint: not yet peer-reviewed. June 2023.
@@
Propose a novel system, chorus, whose flexible architecture enables the synthesis of multiple data discovery tasks and deploying risk mitigations
@@
Introduce novel mitigations, including nearest-neighbor matching and anchoring, to reduce foundation-model risks specific to this domain;
@@ Discussion
This results in a high degree of flexibility. Novel tasks can be specified in natural text, without need for expensive data collection—shown with the example prompts in Figure 5. Another advantage of this approach is a unified architecture: tasks can utilize the overall context and previous outputs. For example, in Figure 1 the the table class of ElectricVehicle can help with deducing the outputs of Make and Model in the next task.
@@
Section 2 defines the three tasks investigated in this paper. Section 3 describes the architecture of chorus and key approaches. We evaluate the performance of chorus in Section 4’s experiments and offer a discussion of those results in Section 5. This includes a discussion of promising directions in Section 5.1. Finally, we place this work within the literature in Section 6, discussing related works.
@@ 2.1. Tasks
In addition to the data collection, we are also given a reference ontology of table classes , and a reference ontology of column types . For example, the DBPedia.org types for the table classes include https://dbpedia.org/ontology/Lake, https://dbpedia.org/ontology/Actor and https://dbpedia.org/ontology/Continent and column types include https://dbpedia.org/ontology/areaTotal and https://dbpedia.org/ontology/birthDate.
We consider three tasks of interest to perform over this data collection.
@@
For each table , determine its appropriate class , such that every row represents an instance of the type. See (Koleva et al., 2021) for more information.
@@
Assume an execution log , which maps many where . Given two tables and , with columns and respectively, the join-column prediction task is to suggest a pair of columns such that the equality condition , which can be used to join the the tables, matches with the choice in the execution log . For more discussion, see (Yan and He, 2020).
@@
2.2. Foundation Models
We discuss some properties of foundation models relevant to this project. For an overview of foundation models and their capabilities, we recommend this comprehensive treatment (Bubeck et al., 2023). For the following sections, we examine the capabilities of the GPT-3.5 model (Ouyang et al., 2022).
Relational data We find that foundation models have the capability of parsing relational data. Consider for example the input table in Figure 1. We serialize the table into a comma-separated values (csv) format. Inputting this into a foundation model, in this case GPT-3.5, easily shows that the model can reason about the relational structure of the data, as seen in Figure 2.
From the English utterances ”header row” and ”second column of the third data row”, the model is able to reason about the provided table and output the intended values. This requires understanding of schemas, relations, tuples, attributes and values. Thus all the basic blocks of the relational model are present in the model.
@@ 2.2. Foundation Models
As earlier, we demonstrate with an example in Figure 3.
This shows that the model encodes information about popular ontology classes and properties. Note that this information is not necessarily complete nor correct. We emphasize the current generation of foundation models does not have access to data lookup abilities, despite them occasionally generating output claiming to have done precisely that (Court, 2023).
@@ 3. Approach
For each task instance, Chorus generates a prompt by concatenating six inputs: a combination of instructions, context, task-specific knowledge, data samples and metadata to form the model input. These are colored-coded in Figure 5. This natural langauge input is then fed to the foundation model. The output is then subject to post-processing: checks of parsability and feasibility are conducted. If these pass, the output is extracted. Otherwise, we activate a mitigation we denote anchoring to repair the error and prevent its propagation.
@@
AcademicJournal, AdministrativeRegion, Airline, Airport, Animal, BaseballPlayer, Bird, Book, Building, City, Company, Country, Cricketer, Currency, Election, FictionalCharacter, Film, GolfPlayer, Hospital, Hotel, Lake, Mammal, Monarch, Mountain, Museum, Newspaper, Novel, Person, Plant, PoliticalParty, RadioStation, Saint, Scientist, Swimmer, TelevisionShow, University, VideoGame, Work, Wrestler.
For example, for a dataset about hospitals, return `https://dbpedia.org/ontology/Hospital `. Begin your answer with’https://dbpedia.org/ontology’
@@
Ford, Transit, 98501, 1FTBW1YKXP,
Hyundai, Ioniq 5, 98027, KM8KRDAF6P,
@@ 3.1. Model Inputs
A description of the specific task (table-class detection, column-type annotation or join-column prediction) is provided to the foundation model in natural language. These are shown in yellow in Figure 5. For example, we translate the formal Definition 2.1 of the first task, table-class detection, into the English sentence “For the following CSV sample, select one DBpedia.org ontology that represents the dataset.” For the third task, join-column prediction, we utilize a code-completion approach. We frame the task as completing Pandas code that performs a join. We choose Pandas because it is a very popular framework, with more than millions of example lines of code on the web. This is the zero-shot prompt setting: the model can be provided with instructions for a novel task and performs them directly.
@@
For the first two tasks, we use the foundation models with additional inputs: this the few-shot prompt setting. The model is given a few demonstrations of task completion, including inputs and outputs. This is shown in Figure 5 as green text.
@@
Foundation models can understand relational data. By serializing the input tables, we can input them into foundation models. For example, consider the example table from Figure 1 in the introduction. Serializing the table allows the foundation model to ingest the data. In this paper, we use the comma-seperated values (csv) format. This gives us the representation in Figure 6.
Because the models have a limited context window size—typically in the few thousands of tokens—tables cannot always be ingested as a whole. Instead, we always serialize a sample of the rows. Intuitively, this is fine because the tasks we consider can be completed without reviewing all the rows, i.e. it is sufficient to consider a few values when determining the type of a column.
@@
Schema information including column names (headers) and keys can be incorporated into the input, above the serialized data sample. We found that foundation models can adaptively infer whether the first column of the input is a header or data row, with no modification of the input required. This is shown in orange in Figure 5. Due to the flexibility of the input format, we add supplemental information to the prompt where available. For example, contextual information about the data source is added, such as “this dataset is from a page titled Washington State Open Government.”
@@
For some tasks, additional information can be used to guide the model. For example, if only certain output classes are desired, these can be listed to the model. An example of such additional constraints for the table-class detection task are shown in Figure 4(a).
@@
3.2. Post-processing
Once the foundation model has run and provided its natural language output, we perform post-processing to parse into a symbolic representation and mitigate common errors.
@@ 3.2. Post-processing
Because the model is not constrained in its outputs, it may not always output a feasible answer. In this setting we impose three constraints: table types must belong to the ontology classes, column types must belong to the ontology properties and joins must be on existing columns. An output may be infeasible if in particular, it is not parsable or if it violates any of the three constraints. To guard against this, we parse the output and check the constraint corresponding to the current task. If the constraint fails, the chorus performs anchoring.
@@
If the constraint is violated, we do not simply reissue the prompt. The reason is that an insidious risk in our setting is hallucination snowballing (Zhang et al., 2023). This recently formalized phenomena describes how once a foundation model makes a spurious generation, subsequent generations are more likely to be erroneous: after a misstep, the system tends to make mistakes it would otherwise be able to avoid. We provide an example in Figure 7: once nonexistent class animalName is suggested, another nonexistent class iucnStatus follows. Because we maintain context across tasks, we are particularly vulnerable to this.
Instead, we resolve this by novel domain-specific mitigation we call anchoring, shown in Figure 7. Chorus maintains the list of embeddings of all feasible answers, e.g. all table classes from DBPedia. The embeddings are computed by running open-source foundation model and extracting the final layer weights. We use for this purpose an alternative foundation model (GPT-3) that allows us access to its weights. Then, for each output that violates the constraints, we conduct the following repair process: (1) We extract the embedding for the incorrect output, using the same process as above. (2) We conduct a nearest-neighbor search on the pre-computed embeddings of the feasible answers. (3) We replace the infeasible output by the nearest-neighbor, transformed into the correct answer format. To the best of our knowledge, our anchoring technique is novel, and only applies to our specific problem where the output is constraints. In contrast, for general-purpose natural language processing the domain is unconstrained.
In the example in Figure 7, we “erase” the model’s history of outputting non-existent class animalName, replacing it with the correct binomial class. After this correction (anchoring), the model is able to generate the correct class for the next column (conservationStatus).
@@ 4. Experiments
We empirically evaluate chorus on the three tasks defined in Section 2.1. For each task, we select a task-specific benchmark and compare with baselines representing the state of the art. Table-class detection 1⃝ in evaluated in Section 4.1, 2⃝ column-type annotation in Section 4.2, and 3⃝ join-column prediction in Section 4.3.
We find that on the table-class detection and join-column prediction tasks, chorus exceeds all tested baselines by a clear margin of and points, respectively. On the column-type annotation task, chorus exceeds all models that are not pretrained on the benchmark dataset (by points) and performs comparably to those which are. We also compare chorus with expert-labels for the table-class detection task, finding exact matches for % of labels and rate the chorus labels as better than the expert-labels when they disagree.
@@
Relevant systems include Tabert (Yin et al., 2020), DoDuo (Suhara et al., 2022), Sato (Zhang et al., 2020), TURL (Deng et al., 2020), TaBBIE (Iida et al., 2021), Auto-suggest (Yan and He, 2020), Trifecta Wrangler (Trifacta, 2023), Paxata, Tableau Prep, and Sherlock. DoDuo is reported to outperform TURL and Sherlock on column-type annotation (Suhara et al., 2022), so we select it for evaluation. Sato and Sherlock are similar, with Sato utilizing additional signals not found in our benchmarks, so we evaluate the better-established Sherlock. TaBBIE can embed tables but is not trained on column-type annotation unlike DoDuo and Tabert, so we avoid it for the column-type annotation task. Tabert is a work similar to DoDuo and TURL, but from the NLP community rather than the data management community, so we also test it too. For join-column prediction, Trifacta Wrangler outperforms Paxata and Tableau Prep (Yan and He, 2020). Auto-Suggest is reported to outperform Trifacta Wrangler, but is a proprietary research project not released publicly. Thus we select Trifacta Wrangler for testing.

### v3

# Chorus: Foundation Models for Unified Data Discovery and Exploration

@@ Abstract.
We apply foundation models to data discovery and exploration tasks. Foundation models are large language models (llms) that show promising performance on a range of diverse tasks unrelated to their training. We show that these models are highly applicable to the data discovery and data exploration domain. When carefully used, they have superior capability on three representative tasks: table-class detection, column-type annotation and join-column prediction. On all three tasks, we show that a foundation-model-based approach outperforms the task-specific models and so the state of the art. Further, our approach often surpasses human-expert task performance. We investigate the fundamental characteristics of this approach including generalizability to several foundation models and the impact of non-determinism on the outputs. All in all, this suggests a future direction in which disparate data management tasks can be unified under foundation models.
@@ 1. Introduction
Despite their promise, serious risks have hampered the reception of foundation models. These include: spurious generation (including “hallucination”) (Holtzman et al., 2020), factual recall limitations (Mallen et al., 2022), bias (Gehman et al., 2020), dataset contamination (Dodge et al., 2021), logical shortcuts (Shi et al., 2023) and fallacies (Lin et al., 2022). Naïve deployment can lead to unanticipated problems: it has already led to legal action (Court, 2023) and recalls by major corporations (Heaven, 2022). These risks are now acknowledged by the creators of these models (Bubeck et al., 2023; Ouyang et al., 2022; Touvron et al., 2023a).
The goal of this paper is to demonstrate the utility of foundation models to the data discovery and exploration while mitigating the aforementioned risks. We select three representative tasks to show the promise of foundation models: 1⃝ table-class detection, 2⃝ column-type annotation and 3⃝ join-column prediction. An outline of our approach is shown in Figure 1. We call our approach chorus. Code for the system and experiments is available at https://github.com/mkyl/CHORUS.
@@
Propose a novel system, chorus, whose flexible architecture enables the synthesis of multiple data discovery tasks and deploying risk mitigations;
@@
Introduce the novel mitigation of anchoring to reduce foundation-model risks specific to this domain;
@@ Discussion
This results in a high degree of flexibility. Novel tasks can be specified in natural text, without need for expensive data collection—task examples, metadata and constraints are all incorporated into the task easily. Another advantage of this approach is a unified architecture: tasks can utilize the overall context and previous outputs. For example, in Figure 1 the table class ElectricVehicle helps with deducing the outputs Make, Model in the next task.
@@
Section 2 defines the three tasks investigated in this paper. Section 3 describes the architecture of chorus and key approaches. We evaluate the performance of chorus in Section 4’s experiments. In that section, we also investigate the fundamental characteristics of this approach. We offer a discussion of those results in Section 5. This includes a discussion of promising future directions. Finally, we place this work within the literature in Section 6, discussing related works.
@@ 2.1. Tasks
In addition to the data collection, we are also given a reference ontology of table classes , and a reference ontology of column types . For example, the DBPedia.org types for the table classes include https://dbpedia.org/ontology/Actor and https://dbpedia.org/ontology/Continent and column types include https://dbpedia.org/ontology/areaTotal and https://dbpedia.org/ontology/birthDate.
We consider three tasks of interest on the data collection:
@@
For each table , determine its appropriate class , such that every row represents an instance of the type. We adopt this definition from (Koleva et al., 2021).
@@
Assume an execution log , a history of user actions including table joins and their join conditions, which maps many where . Given two tables and , with columns and respectively, the join-column prediction task is to suggest a pair of columns such that the equality condition , which can be used to join the the tables, matches with the choice in the execution log . For more discussion, see (Yan and He, 2020).
@@ 3. Approach
For each task instance, Chorus generates a prompt by concatenating six inputs: context, demonstration, data samples, metadata, task-specific knowledge, and prefixes. They form the “Model Inputs” box in Figure 2 and are color-coded so that they match the colored prompt components in Figure 3. This natural language input is then fed to the foundation model. The output is then subject to post-processing: checks of parsability and feasibility are conducted. If these pass, the output is extracted. Otherwise, we activate a mitigation process, called anchoring, in order to repair the error and prevent its propagation.
@@
AcademicJournal, AdministrativeRegion, Airline, Airport, […], University, VideoGame, Work, Wrestler.
For example, for a dataset about hospitals, return `https://dbpedia.org/ontology/Hospital`. Begin your answer with ’https://dbpedia.org/ontology’.
@@ 3.1. Model Inputs
A description of the specific task (table-class detection, column-type annotation or join-column prediction) is provided to the foundation model in natural language. These are shown in yellow in Figure 3. For example, we translate the formal Definition 2.1 of the first task, table-class detection, into the English sentence “For the following CSV sample, select one DBpedia.org ontology that represents the dataset.” For the third task, join-column prediction, we utilize a code-completion approach. We frame the task as code-completing a Pandas fragment that performs a join, with the code to complete shown in Figure 2(c). We choose Pandas because it is a very popular framework, with more than millions of example lines of code on the web. This is the zero-shot prompt setting: the model can be provided with instructions for a novel task and performs them directly.
@@
For the first two tasks, we use the foundation models with task examples as an additional input: this is called the few-shot prompt setting. The model is given a few demonstrations of task completion, including inputs and outputs. This is shown in Figure 3 as green text.
@@
By serializing the input tables, we can input them into foundation models. For example, consider the example table from Figure 1 in the introduction. Serializing the table allows the foundation model to ingest the data. We use the comma-separated values (csv) format, shown in blue in Figure 3.
Because the models have a limited context window size—typically in the few thousands of tokens—tables cannot always be ingested as a whole. Instead, we always serialize a sample of the rows. We find a sample size of five is sufficient. Intuitively, it suffices to consider only a few values to determine column type.
@@
Schema information including column names (headers) and keys can be incorporated into the input, above the serialized data sample. We found that foundation models can adaptively infer whether the first column of the input is a header or data row, with no modification of the input required. This is shown in orange in Figure 3.
@@
For some tasks, additional information can be used to guide the model. For 1⃝ table-class detection, if only certain output classes are desired, these can be listed to the model. The model will take these instructions into account when generating an output but they are not hard constraints. The encoding of such additional constraints for the table-class detection task is shown in Figure 2(a).
@@
3.2. Model Harness
The foundation model is run within a harness that parses outputs into a symbolic representation and mitigates errors.
@@ 3.2. Model Harness
Because the model is not constrained in its outputs, it may not always output a feasible answer. In this setting we impose three constraints: table types must belong to the ontology classes, column types must belong to the ontology properties and joins must be on existing columns. An output is infeasible if, in particular, it is not parsable or if it violates any of the three constraints. If this occurs, chorus performs anchoring.
@@
If the constraints are violated, we do not simply move on to the next task. The risk is of hallucination snowballing (Zhang et al., 2023): once a foundation model makes a single spurious generation, subsequent outputs are more likely to also be wrong. The model will make mistakes it would otherwise be able to avoid. For example, in Figure 4(a): once nonexistent class iucnStatus is suggested, another nonexistent class animalName follows. Because we maintain context across tasks, we are particularly vulnerable to this.
We call the novel domain-specific mitigation we deploy anchoring, shown in Figure 4(b). Chorus ends the conversation when an error is detected. It then initiates a new conversation, feeding the llm with a false history in which the llm did not hallucinate. This is possible because the conversational llm takes as input the full history text, which we can retroactively modify. We insert artificially an existing class from the ontology (e.g. the nearest neighbor in the embedding space to the non-existing class). Fed with this cleaner input, the model is able to directly provide the correct answer.
@@ 4. Experiments
We empirically evaluate chorus on the three tasks defined in Section 2.1. For each task, we select a task-specific benchmark and compare with baselines representing the state of the art. Table-class detection 1⃝ is evaluated in Section 4.1, 2⃝ column-type annotation in Section 4.2, and 3⃝ join-column prediction in Section 4.3. The code for the experiments in available at https://github.com/mkyl/CHORUS.
@@
We considered the following state-of-the-art systems for data exploration: relevant systems include Tabert (Yin et al., 2020), DoDuo (Suhara et al., 2022), Sato (Zhang et al., 2020), TURL (Deng et al., 2020), TaBBIE (Iida et al., 2021), Auto-suggest (Yan and He, 2020), Trifecta Wrangler (Trifacta, 2023), Paxata, Tableau Prep, and Sherlock. DoDuo outperforms TURL and Sherlock on column-type annotation (Suhara et al., 2022), so we select it for evaluation. Sato and Sherlock are similar, with Sato utilizing additional signals not found in our benchmarks, so we evaluate the better-established Sherlock. TaBBIE can embed tables but is not trained on column-type annotation unlike DoDuo and Tabert, so we avoid it for the column-type annotation task. Tabert is a work similar to DoDuo and TURL, but from the NLP community rather than the data management community, so we also test it too. For join-column prediction, Trifacta Wrangler outperforms Paxata and Tableau Prep (Yan and He, 2020). Auto-Suggest is reported to outperform Trifacta Wrangler, but is a proprietary research project not released publicly. Thus we select Trifacta Wrangler for testing.